Diffusion Morning Report 11/17/05

Ted Berg tberg at snf.stanford.edu
Thu Nov 17 06:34:27 PST 2005


Good Morning All,
    Looking Ok this morning. Yesterday was quite busy. Nitride pump was 
changed and Epi was taken down for tube change.  Epi is being re-
assembled . Nitride is fine again. Pump lines for Thermco 2 should be 
installed this week and that should get us close to the end of the 
facilities
part of the install.  SiGe Germane is in leak check. Everything else 
looks ok for now.

  	Date: 	11/17/2005 	THICKNESS-Change at
Tube Name/Number 	Status 	Comments 	 
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	Up 	  	 
Tylan 8-SiGe 	Up 	  	10.42/30
Tylan 9-Poly 	Up 	  	28.16/30
Tylan 10-Nitride 	Up 	Pump changed yesterday 	7.0/30 
Tylan 11-BPSG-LTO 	Up 	  	11.22/30
Tylan 12-FGA 	Up 	  	 
  	  	  	 
Thermco 1 	yellow 	problem with front zone temp 	 
Thermco 2 	Up 	  	 
Thermco Poly 	Up 	  	 
Thermco Nitride 	Down 	Testing 	 
  	  	  	 
Burn Box 	Up 	  	 
ZYGO 	up 	  	.
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	B on line 390 	 changeover working well 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20051117/b37bd1cd/attachment.html>


More information about the diffusion mailing list